OpenStudy (anonymous):

differentiate 4x+6/x^2

OpenStudy (anonymous):

4x-12/x^3?

OpenStudy (jhannybean):

Wait a minute.... If your problem \[4x+\frac{6}{x^2}\] oris it \[\frac{4x+6}{x^2}\]

OpenStudy (anonymous):

the first one

OpenStudy (jhannybean):

Alright, then that's different.

OpenStudy (jhannybean):

Take the derivative with respect to each part. \[\frac{d}{dx}(4x) +6\frac{d}{dx}\left(\frac{1}{x^2}\right)\]

OpenStudy (jhannybean):

The derivative of x is just 1, so \(\frac{d}{dx}(4x)=4\) \[6\cdot \frac{d}{dx}\left(\frac{1}{x^2}\right)=6\cdot \frac{d}{dx}(x^{-2})\] So what is the derivative of x\(^{-2}\)?

OpenStudy (jhannybean):

And yeah, I hadnt bothered solving it, so I was working it out here. But you are correct.